Two charged particles exert an electrostatic force of 7 N on each other. What will the magnitude of the electrostatic force be if the distance between the two charges is reduced to one-fourth of the original distance?
Answer #1

We know force between two charge particles

F=k Q 1* Q 2/ R 2 =7 N ..1

according to the question R'=R/4

F'=k Q 1*Q 2 /R'2

so F'=k Q 1 *Q 2 /(R/4)2

F'=16*k Q 1*Q 2 /R 2 ..2

using equation 1 in 2

F'=16*7

F'= 112 N Answer.